# Spatial Scale of Agglomeration and Dispersion Number, Spacing, and the Spatial Extent of Cities

### v7

@@ Appendix F Derivations
This section collects omitted derivations. Below, the expression $\mathbf{F}_{x}$ denotes the Jacobian matrix of a vector-valued function $\bm{f}(\bm{x})$ with respect to $\bm{x}$, that is, $\mathbf{F}_{x}=[\frac{\partial f_{i}}{\partial x_{j}}]$. For example, $\mathbf{V}_{x}\equiv[\frac{\partial v_{i}}{\partial x_{j}}]$, $\tilde{\mathbf{V}}_{w}\equiv[\frac{\partial\tilde{v}_{i}}{\partial w_{j}}]$, and $\mathbf{W}_{x}\equiv[\frac{\partial w_{i}}{\partial x_{j}}]$. Throughout, $\bar{v}$, $\bar{w}$, $\bar{e}$ and so on represent $v_{i}$, $w_{i}$, $e_{i}$ evaluated at $\bar{\bm{x}}$. $\mathbf{D}$ denotes the row-normalized proximity matrix.
@@ F.1.1   The benefit matrix
In spatial models, the payoff function (i.e., indirect utility function) $\bm{v}$ often reduce to the following implicit form:
@@
The condition Eq. F.2 represents, e.g., the general equilibrium conditions for a given $\bm{x}$ that defines endogenous variable $\bm{w}$ (e.g., wages) other than $\bm{x}$ as an implicit function of $\bm{x}$. For $\bm{v}(\bm{x})$ to be well-defined, Eq. F.2 must admit a unique solution of $\bm{w}$ for all $\bm{x}\in\mathcal{X}$. Suppose $\bm{s}$ and $\bm{v}$ are continuously differentiable. Then, we have
@@
where $\mathbf{W}_{x}(\bm{x})$ is obtained by applying the implicit function theorem to Eq. F.2.
Under Assumption C, all relevant matrices commute at $\bm{x}=\bar{\bm{x}}$ because they are real, symmetric, and circulant at $\bar{\bm{x}}$. Thus, $\mathbf{V}_{x}=\mathbf{S}_{w}^{-1}(\mathbf{S}_{w}\tilde{\mathbf{V}}_{x}-\tilde{\mathbf{V}}_{w}\mathbf{S}_{x})$ at $\bar{\bm{x}}$.
@@
where regional expenditure is $e_{i}=e(w_{i},x_{i})$ with some nonnegative function $e$ and $\mathbf{M}=[m_{ij}]$ is the expenditure share matrix whose typical element being
@@
Suppose Assumption C. Suppose $\bm{x}=\bar{\bm{x}}$ and let $\bar{w}$ be the uniform level of $\{w_{i}\}$ at $\bar{\bm{x}}$. Then, we have $\mathbf{M}=\mathbf{D}$ at $\bm{x}=\bar{\bm{x}}$. Suppose the scalars $\epsilon_{x}$ and $\epsilon_{w}$ are chosen to satisfy $\mathbf{E}_{x}=\epsilon_{x}\bar{w}\mathbf{I}$ and $\mathbf{E}_{w}=\epsilon_{w}\bar{x}\mathbf{I}$ at $\bar{\bm{x}}$. Let $\bar{e}=e(\bar{w},\bar{x})$ and $\zeta\equiv\frac{\bar{e}}{\bar{w}\bar{x}}$. We see that
@@ F.1.2   The payoff elasticity matrix with respect to local characteristics
In Eq. E.15, $\mathbf{X}=[\frac{\partial x_{i}(\bar{\bm{a}})}{\partial a_{i}}]=\mathbf{X}_{a}$ acts as $\hat{\mathbf{X}}\equiv-\mathbf{V}_{x}^{-1}\mathbf{V}_{a}$ for $\bm{z}$ such that $\bm{z}^{\top}\bm{1}=\bm{0}$. Thus, $\mathbf{V}_{a}$ is of interest.
For purely local characteristics (Example 3), since $v_{i}(\bm{x},\bm{a})=a_{i}v_{i}(\bm{x})$, it follows that $\mathbf{V}_{a}=\operatornamewithlimits{diag}[\bm{v}(\bm{x})]$. At $\bar{\bm{x}}$, we have $\mathbf{V}_{a}=\bar{v}\mathbf{I}$. Thus, $\hat{\mathbf{X}}=-\bar{v}\mathbf{V}_{x}^{-1}$.
For regional characteristics that affect trade flows, the payoff function and the market equilibrium condition are, respectively, modified to $\bm{v}(\bm{x},\bm{a})=\tilde{\bm{v}}(\bm{x},\bm{w},\bm{a})$ and $\bm{s}(\bm{x},\bm{w},\bm{a})=\bm{0}$. By applying the implicit function theorem, we see $\mathbf{V}_{a}=\tilde{\mathbf{V}}_{a}+\tilde{\mathbf{V}}_{w}\mathbf{W}_{a}=\tilde{\mathbf{V}}_{a}-\tilde{\mathbf{V}}_{w}\mathbf{S}_{w}^{-1}\mathbf{S}_{a}$. As all matrices commute at $\bar{\bm{x}}$ under Assumption C, it is equivalent to consider
@@
So, $\mathbf{S}_{a}=-\left(\operatornamewithlimits{diag}[\mathbf{M}\bm{e}]-\mathbf{M}\operatornamewithlimits{diag}[\bm{e}]\mathbf{M}^{\top}\right)\operatornamewithlimits{diag}[\bm{a}]^{-1}=-\frac{\bar{e}}{\bar{a}}\left(\mathbf{I}-\mathbf{D}^{2}\right)$. See Section F.2.2. ∎
@@
where $e$ maps the tuple $(w_{j},x_{j},a_{j})$ to the regional expenditure. Then, we have $\mathbf{S}_{a}=-\mathbf{M}\mathbf{E}_{a}$, or $\mathbf{S}_{a}=-\epsilon_{a}\mathbf{D}$ at $\bar{\bm{x}}$ where $\epsilon_{a}=\frac{\partial e(\bar{x},\bar{w},\bar{a})}{\partial a_{i}}$. See Section F.2.1. ∎
@@ F.2   Model-specific derivations
We provide omitted derivations of the gain functions $\Omega$, as defined in Section A.2, for the examples in the main text. For derivations for other models mentioned in Sections 2.5 and 2.6, see Akamatsu-et-al-DP2017, an earlier draft of the current paper.
@@ F.2.1   Krugman-JPE1991 model
There are two types of workers, mobile and immobile, and their total masses are $1$ and $L$, respectively. $\bm{x}\equiv(x_{i})_{i\in\mathcal{I}}$ is the distribution of mobile workers. Each worker supplies one unit of labor inelastically.
@@
In the M-sector, to produce $q$ units, a firm requires $\alpha+\beta q$ units of mobile labor. Profit maximization of firms yields the price of differentiated goods produced in region $i$ and exported to $j$ as $p_{ij}=\frac{\sigma\beta}{\sigma-1}w_{i}\tau_{ij}$, which in turn determines gravity trade flow from $j$ to $i$. That is, when $X_{ij}$ denotes the price of M-sector goods produced in region $i$ and sold in region $j$, $X_{ij}=m_{ij}e_{j}$ where the share $m_{ij}\in(0,1)$ is defined by Eq. F.6 with $\phi_{ij}\equiv\tau_{ij}^{1-\sigma}$. The proximity matrix is thus $[\phi_{ij}]=[\tau_{ij}^{1-\sigma}]$.
Given $\bm{x}$, we determine the market wage $\bm{w}\equiv(w_{i})_{i\in\mathcal{I}}$ by the M-sector product market-clearing, zero-profit, and mobile labor market-clearing conditions. These conditions are summarized by the trade balance $w_{i}x_{i}=\sum_{j\in\mathcal{I}}X_{ij}$, or Eq. F.5 with $e(x_{i},w_{i})=\mu(w_{i}x_{i}+l_{i})$. By adding up Eq. F.5 for the Krugman model, we see $\sum_{i\in\mathcal{I}}w_{i}x_{i}=\frac{\mu}{1-\mu}L$, which constrains the total income of mobile workers at any configuration $\bm{x}$. The existence and uniqueness of the solution for Eq. F.5 follow from standard arguments (e.g., Facchinei-Pang-Book2007). Given the solution $\bm{w}(\bm{x})$ of Eq. F.5, we have the indirect utility of mobile workers, which is given by $v_{i}=\Delta_{i}^{\frac{\mu}{\sigma-1}}w_{i}$, where $\Delta_{i}\equiv\sum_{k\in\mathcal{I}}x_{k}w_{k}^{1-\sigma}d_{ki}$.
@@
where Eq. F.4 and Eq. F.8 give $\mathbf{W}_{x}$. By plugging $\delta=\frac{\mu(\bar{w}\bar{x}+l)}{\bar{w}\bar{x}}=1$ and $\epsilon_{x}=\epsilon_{w}=\mu$ to Eq. F.8,
@@
From Eq. F.18 we have $\mathbf{V}=\Omega^{\flat}(\mathbf{D})^{-1}\Omega^{\sharp}(\mathbf{D})$, where we define
@@
Using the Krugman model as an example, we discuss how economic forces in a model are embedded in $\Omega$ and $\Omega^{\sharp}$. We recall that positive (negative) terms in $\Omega$ represent agglomeration (dispersion) forces. In Eq. F.18, (a) corresponds to the elasticity of price index with respect to agents’ spatial distribution $\bm{x}$, (b) to the elasticity of payoff with respect to nominal wage $\bm{w}$, (c) to the elasticity of wage with respect to agents’ spatial distribution. Here, (a) and the second term in (b) corresponds to the so-called cost-of-living effect through price index; (a) is positive, i.e., it is an agglomeration force, as the price index in a region becomes lower when more agents (firms) locate geographically close regions; the second term in (b) (i.e., $-\mu\Theta$) is negative because higher wage in a region implies higher goods prices in its nearby regions. Also, (b) as a whole is positive, meaning that the payoff of a region is increasing in wages even with the negative effect through price index. The last component (c) includes both positive and negative terms; in its numerator, the first term ($\mu\Theta$) is demand linkage where firms’ profits rise when they are close to regions with high total income, and the second term ($-\Theta^{2}$) is the market-crowding effect due to competition between firms. The sign of (c) is $\Theta$-dependent; for example, it is negative when $\Theta$ is high ($\phi$ is low) and positive otherwise. The denominator of (c) represents the general equilibrium effects through the so-called short-run equilibrium condition under given $\bm{x}$, i.e., Eq. F.5. As $\Omega^{\sharp}$ is obtained by combining these components and collecting terms according to the order of $\Theta$, these economic forces affect both the first- and second-order coefficients of $\Omega^{\sharp}$. Concretely, in Eq. F.19, $\frac{\mu}{\sigma-1}\Theta$ comes from (a), $\frac{\mu}{\sigma}\Theta$ comes from (b) $\times$ (c), $-\frac{1}{\sigma}\Theta^{2}$ comes from (b) $\times$ (c), and $-\frac{\mu^{2}}{\sigma-1}\Theta^{2}=-\left(\frac{\mu^{2}}{\sigma}+\frac{\mu^{2}}{\sigma(\sigma-1)}\right)\Theta^{2}$ comes from all three components while its leading term $-\frac{\mu^{2}}{\sigma}\Theta^{2}$ comes from (b) $\times$ (c). Thus, by considering $\Omega^{\sharp}$ for a model, one can examine the net effect of all economic forces in the model at once, and the net effect is decomposed according to its spatial scale (i.e., the order of $\Theta$). ∎
@@
In Fig. C.3, we set $\mu=0.5$, $\sigma=10$, and $L=8$. ∎
Remark 7.
To obtain $\Omega^{\natural}$ for $\bm{l}=(l_{i})_{i\in\mathcal{I}}$, we evaluate $\mathbf{V}_{l}=-\tilde{\mathbf{V}}_{w}\mathbf{S}_{w}^{-1}\mathbf{S}_{l}$ as $\mathbf{A}=\frac{l}{\bar{v}}\mathbf{V}_{l}$. From Example 6, $\mathbf{S}_{l}=-\mu\mathbf{D}$. Also, $\tilde{\mathbf{V}}_{w}=\bar{v}\frac{\partial}{\partial\bm{w}}\log\bm{v}(\bar{\bm{x}})=\frac{\bar{v}}{\bar{w}}(\mathbf{I}-\mu\mathbf{D})$ and $\tilde{\mathbf{V}}_{l}=\bm{0}$. Thus,

### v8

@@ Appendix F Derivations
This section collects omitted derivations. The expression $\mathbf{F}_{x}$ denotes the Jacobian matrix of a vector-valued function $\bm{f}(\bm{x})$ with respect to $\bm{x}$, that is, $\mathbf{F}_{x}=[\frac{\partial f_{i}}{\partial x_{j}}]$. For example, $\mathbf{V}_{x}\equiv[\frac{\partial v_{i}}{\partial x_{j}}]$, $\tilde{\mathbf{V}}_{w}\equiv[\frac{\partial\tilde{v}_{i}}{\partial w_{j}}]$, and $\mathbf{W}_{x}\equiv[\frac{\partial w_{i}}{\partial x_{j}}]$. Throughout, $\bar{v}$, $\bar{w}$, $\bar{e}$ and so on represent $v_{i}$, $w_{i}$, $e_{i}$ evaluated at $\bar{\bm{x}}$. $\mathbf{D}$ denotes the row-normalized proximity matrix.
@@ F.1.1  The benefit matrix
The indirect utility function $\bm{v}$ of a spatial model often reduce to the following implicit form:
@@
The condition Eq.˜F.2 represents, e.g., the general equilibrium conditions for a given $\bm{x}$ that defines endogenous variable $\bm{w}$ (e.g., wages) other than $\bm{x}$ as an implicit function of $\bm{x}$. We assume that Eq.˜F.2 admits a unique solution of $\bm{w}$ at each $\bm{x}$ for $\bm{v}(\bm{x})$ to be well-defined. Suppose $\bm{s}$ and $\bm{v}$ are continuously differentiable. Then, we have
@@
where $\mathbf{W}_{x}(\bm{x})$ is obtained by applying the implicit function theorem to Eq.˜F.2.
Under Assumption˜C, all relevant matrices commute at $\bm{x}=\bar{\bm{x}}$ because they are real, symmetric, and circulant at $\bar{\bm{x}}$. Thus, $\mathbf{V}_{x}=\mathbf{S}_{w}^{-1}(\mathbf{S}_{w}\tilde{\mathbf{V}}_{x}-\tilde{\mathbf{V}}_{w}\mathbf{S}_{x})$ at $\bar{\bm{x}}$.
@@
where regional expenditure is $e_{i}=e(w_{i},x_{i})$ with some nonnegative function $e$ and $\mathbf{M}=[m_{ij}]$ is the expenditure share matrix. For example, in the Krugman and Helpman models,
@@
Suppose Assumption˜C. Suppose $\bm{x}=\bar{\bm{x}}$ and let $\bar{w}$ be the uniform level of $\{w_{i}\}$ at $\bar{\bm{x}}$. Then, we have $\mathbf{M}=\mathbf{D}$ at $\bm{x}=\bar{\bm{x}}$. Suppose the scalars $\epsilon_{x}$ and $\epsilon_{w}$ are chosen to satisfy $\mathbf{E}_{x}=\epsilon_{x}\bar{w}\mathbf{I}$ and $\mathbf{E}_{w}=\epsilon_{w}\bar{x}\mathbf{I}$ at $\bar{\bm{x}}$. Let $\bar{e}=e(\bar{w},\bar{x})$ and $\zeta\equiv\frac{\bar{e}}{\bar{w}\bar{x}}$. We see that
@@ F.1.2  The payoff elasticity matrix with respect to local characteristics
In Eq.˜E.12, $\mathbf{X}=[\frac{\partial x_{i}(\bar{\bm{a}})}{\partial a_{i}}]=\mathbf{X}_{a}$ acts as $\hat{\mathbf{X}}\equiv-\mathbf{V}_{x}^{-1}\mathbf{V}_{a}$ for $\bm{z}$ such that $\bm{z}^{\top}\bm{1}=\bm{0}$. Thus, $\mathbf{V}_{a}$ is of interest.
For purely local characteristics (Example˜5), since $v_{i}(\bm{x},\bm{a})=a_{i}v_{i}(\bm{x})$, it follows that $\mathbf{V}_{a}=\operatornamewithlimits{diag}[\bm{v}(\bm{x})]$. At $\bar{\bm{x}}$, we have $\mathbf{V}_{a}=\bar{v}\mathbf{I}$. Thus, $\hat{\mathbf{X}}=-\bar{v}\mathbf{V}_{x}^{-1}$.
For regional characteristics that affect trade flows, the payoff function and the market equilibrium condition are, respectively, modified to $\bm{v}(\bm{x},\bm{a})=\tilde{\bm{v}}(\bm{x},\bm{w},\bm{a})$ and $\bm{s}(\bm{x},\bm{w},\bm{a})=\bm{0}$. By applying the implicit function theorem, we see $\mathbf{V}_{a}=\tilde{\mathbf{V}}_{a}+\tilde{\mathbf{V}}_{w}\mathbf{W}_{a}=\tilde{\mathbf{V}}_{a}-\tilde{\mathbf{V}}_{w}\mathbf{S}_{w}^{-1}\mathbf{S}_{a}$. As all matrices commute at $\bar{\bm{x}}$ under Assumption˜C, it is equivalent to consider
@@
Thus, $\mathbf{S}_{a}=-\left(\operatornamewithlimits{diag}[\mathbf{M}\bm{e}]-\mathbf{M}\operatornamewithlimits{diag}[\bm{e}]\mathbf{M}^{\top}\right)\operatornamewithlimits{diag}[\bm{a}]^{-1}=-\frac{\bar{e}}{\bar{a}}\left(\mathbf{I}-\mathbf{D}^{2}\right)$. See Section˜F.2.2. ∎
@@
where $e$ maps the tuple $(w_{j},x_{j},a_{j})$ to the regional expenditure. Then, we have $\mathbf{S}_{a}=-\mathbf{M}\mathbf{E}_{a}$, or $\mathbf{S}_{a}=-\epsilon_{a}\mathbf{D}$ at $\bar{\bm{x}}$ where $\epsilon_{a}=\frac{\partial e(\bar{x},\bar{w},\bar{a})}{\partial a_{i}}$. See Section˜F.2.1. ∎
@@ F.2  Model-specific derivations
We provide omitted derivations of the gain functions $\Omega$, as defined in Section˜A.3, for the examples in the main text. For derivations for other models mentioned in Sections˜2.4 and 2.5, see Akamatsu-et-al-DP2017, an earlier draft of the current paper.
@@ F.2.1  Krugman-JPE1991 model
There are two types of workers, mobile and immobile, and their total masses are $1$ and LL, respectively. $\bm{x}\equiv(x_{i})_{i\in\mathcal{I}}$ is the distribution of mobile workers. Each worker supplies one unit of labor inelastically.
@@
In the M-sector, to produce $q$ units, a firm requires $\alpha+\beta q$ units of mobile labor. Profit maximization of firms yields the price of differentiated goods produced in region $i$ and exported to $j$ as $p_{ij}=\frac{\sigma\beta}{\sigma-1}w_{i}\tau_{ij}$, which in turn determines gravity trade flow from $j$ to $i$. That is, when $X_{ij}$ denotes the price of M-sector goods produced in region $i$ and sold in region $j$, $X_{ij}=m_{ij}e_{j}$ where the share $m_{ij}\in(0,1)$ is defined by Eq.˜F.6 with $\phi_{ij}\equiv\tau_{ij}^{1-\sigma}$. The proximity matrix is thus $\phi_{ij}]=[\tau_{ij}^{1-\sigma}]$.
Given $\bm{x}$, we determine the market wage $\bm{w}\equiv(w_{i})_{i\in\mathcal{I}}$ by the M-sector product market-clearing, zero-profit, and mobile labor market-clearing conditions. These conditions are summarized by the trade balance $w_{i}x_{i}=\sum_{j\in\mathcal{I}}X_{ij}$, or Eq.˜F.5 with $e(x_{i},w_{i})=\mu(w_{i}x_{i}+l_{i})$. By adding up Eq.˜F.5 for the Krugman model, we see $\sum_{i\in\mathcal{I}}w_{i}x_{i}=\frac{\mu}{1-\mu}L$, which constrains the total income of mobile workers at any configuration $\bm{x}$. The existence and uniqueness of the solution for Eq.˜F.5 follow from standard arguments (e.g., Facchinei-Pang-Book2007). Given the solution $\bm{w}(\bm{x})$ of Eq.˜F.5, we have the indirect utility of mobile workers, which is given by $v_{i}=\Delta_{i}^{\frac{\mu}{\sigma-1}}w_{i}$, where $\Delta_{i}\equiv\sum_{k\in\mathcal{I}}x_{k}w_{k}^{1-\sigma}d_{ki}$.
@@
where Eqs.˜F.4 and F.8 give $\mathbf{W}_{x}$. By plugging $\delta=\frac{\mu(\bar{w}\bar{x}+l)}{\bar{w}\bar{x}}=1$ and $\epsilon_{x}=\epsilon_{w}=\mu$ to Eq.˜F.8,
@@
From Eq.˜F.18 we have $\mathbf{V}=\Omega^{\flat}(\mathbf{D})^{-1}\Omega^{\sharp}(\mathbf{D})$, where we define
@@
Using the Krugman model as an example, we discuss how economic forces in a model are embedded in $\Omega$. We recall that positive (negative) terms in $\Omega$ represent agglomeration (dispersion) forces. In Eq.˜F.18, (a) corresponds to the elasticity of price index with respect to agents’ spatial distribution $\bm{x}$, (b) to the elasticity of payoff with respect to nominal wage $\bm{w}$, (c) to the elasticity of wage with respect to agents’ spatial distribution. Here, (a) and the second term in (b) corresponds to the so-called cost-of-living effect through price index; (a) is positive, i.e., it is an agglomeration force, as the price index in a region becomes lower when more agents (firms) locate geographically close regions; the second term in (b) (i.e., $\mu\Theta$) is negative because higher wage in a region implies higher goods prices in its nearby regions. Also, (b) as a whole is positive, meaning that the payoff of a region is increasing in wages even with the negative effect through price index. The last component (c) includes both positive and negative terms; in its numerator, the first term ($\mu\Theta$) is demand linkage where firms’ profits rise when they are close to regions with high total income, and the second term ($\Theta^{2}$) is the market-crowding effect due to competition between firms. The sign of (c) is $\Theta$-dependent; for example, it is negative when $\Theta$ is high ($\phi$ is low) and positive otherwise. The denominator of (c) represents the general equilibrium effects through the so-called short-run equilibrium condition under given $\bm{x}$, i.e., Eq.˜F.5. As $\Omega^{\sharp}$ is obtained by combining these components and collecting terms according to the order of $\Theta$, these economic forces affect both the first- and second-order coefficients of $\Omega^{\sharp}$. Concretely, in Eq.˜F.19, $\frac{\mu}{\sigma-1}\Theta$ comes from (a), $\frac{\mu}{\sigma}\Theta$ comes from (b) $\times$ (c), $\frac{1}{\sigma}\Theta^{2}$ comes from (b) $\times$ (c), and $\frac{\mu^{2}}{\sigma-1}\Theta^{2}=-\left(\frac{\mu^{2}}{\sigma}+\frac{\mu^{2}}{\sigma(\sigma-1)}\right)\Theta^{2}$ comes from all three components while its leading term $\frac{\mu^{2}}{\sigma}\Theta^{2}$ comes from (b) $\times$ (c). Thus, by considering $\Omega^{\sharp}$ for a model, one can examine the net effect of all economic forces in the model at once, and the net effect is decomposed according to its spatial scale (i.e., the order of $\Theta$). ∎
@@
To obtain $\Omega^{\natural}$ for $\bm{l}=(l_{i})_{i\in\mathcal{I}}$, we evaluate $\mathbf{V}_{l}=-\tilde{\mathbf{V}}_{w}\mathbf{S}_{w}^{-1}\mathbf{S}_{l}$ as $\mathbf{A}=\frac{l}{\bar{v}}\mathbf{V}_{l}$. From Example˜8, $\mathbf{S}_{l}=-\mu\mathbf{D}$. Also, $\tilde{\mathbf{V}}_{w}=\bar{v}\frac{\partial}{\partial\bm{w}}\log\bm{v}(\bar{\bm{x}})=\frac{\bar{v}}{\bar{w}}(\mathbf{I}-\mu\mathbf{D})$ and $\tilde{\mathbf{V}}_{l}=\bm{0}$. Thus,